AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  DESPEC, Page 1 of 37  ELOG logo
ID Date Author Subject Text Attachments
  741   Thu Nov 6 12:58:09 2025 NH207Bi source test - data takingRestart AIDA

aida06 took three attempts to boot today
 6x 
  740   Wed Nov 5 13:23:38 2025 JB, MP207Bi source test - FWHM tests14:10 started up the DAQ. HV and TEMPS OK.

Peak threshold with 1V pulser measured at
 8x 
  739   Tue Nov 4 17:29:28 2025 MPBi207 source data, 700keVAnydesk has crashed. I am not able to set
it up again. I switch off bias and fees

  
  738   Tue Nov 4 16:34:47 2025 MP, NHBi207 source data, 700keVSome FEES were not acquiring data for some
reason we could not understand. We lowered
the thresholds to 200 keV, then rose them
  
  737   Tue Nov 4 14:57:19 2025 MPBi207 source data, 700keV- Bias and basic checks: the pulser FWHM
appears to be ~170, which is the same we
have obtained at the end of last week.
 Screenshot_from_2025-11-04_15-49-20.pngScreenshot_from_2025-11-04_15-33-11.pngScreenshot_from_2025-11-04_15-28-42.pngScreenshot_from_2025-11-04_15-28-35.png 
  736   Thu Oct 30 16:43:36 2025 MP, NHSetup and data taking with 207BiData taking and DAQ stopped now (17.40)




  
  735   Thu Oct 30 10:06:21 2025 MP, NHSetup and data taking with 207Bi- Initialised the Merger and Tape server

- loading of settings: EXPERIMENTS/AIDA/2025Oct29-16.11.06

- aida16 very noisy (attachments
 Screenshot_from_2025-10-30_11-07-55.pngScreenshot_from_2025-10-30_11-07-39.pngScreenshot_from_2025-10-30_11-07-29.pngScreenshot_from_2025-10-30_11-08-07.pngScreenshot_from_2025-10-30_11-10-06.png 
  734   Wed Oct 29 14:11:01 2025 JB, MP207Bi source test - data taking15:10 MP entered and check the water pressure
OK.
15:10 DSSSDs biased - OK.
 7x 
  733   Wed Oct 29 13:39:13 2025 TDAIDA FEE64 adaptor PCB to AIDA DSSSD cabling
The cables form the AIDA FEE64 adaptor PCB
to the MSL type BB18 DSSSDs are constructed
  
  732   Wed Oct 29 13:01:08 2025 NHUpdate of AnyDesk AnyDesk was updated to the latest version
that supports CentOS 7 (EOL OS)
  
  731   Wed Oct 29 12:56:29 2025 TDFurther setup for 207Bi testRemember that 'ADC calibration'
refers to the setup of the fast serial output
of the AD9222 Octal, 50MSPS, 14 bit ADC which
  
  730   Wed Oct 29 08:12:35 2025 MPFurther setup for 207Bi testInspection of side FEEs (2,6,4,8) and possible
grounding improvements:

- all screws and nuts for copper
 IMG_5959.jpegIMG_5958.jpegIMG_5957.jpegScreenshot_from_2025-10-29_09-23-32.pngScreenshot_from_2025-10-29_09-32-52.png 
  729   Tue Oct 28 12:43:38 2025 JB, GB, MP, AMFurther setup of tests for 207Bi test- Ch 2 and 3 of CAEN HV module changed to
negative polarity, Ch 0,1 still with positive
polarity
 14x 
  728   Mon Oct 27 18:08:11 2025 TDChanging ground setup from 2510202519.06 Found HV OFF, DAQ GOing

      DAQ STOP
  
  727   Mon Oct 27 11:55:00 2025 JB, MP, GB, NHChanging ground setup from 2510202512:55 We are going to try the follow changes
to the AIDA scheme of 241012025.
 9x 
  726   Sun Oct 26 16:22:38 2025 TD[How To] Check ASIC settings
tcsh
cd /MIDAS/DB/EXPERIMENTS/AIDA/<ASIC settings
  
  725   Fri Oct 24 14:33:37 2025 JB, GB, MP, AMSnout grounding, bias and test input layout as of 24.10.2025   AIDA_scheme_241020525(2).pdf 
  724   Fri Oct 24 12:29:02 2025 JB, GB, MP, AM, NH207Bi source threshold tests> > > 
> > > We have dismounted the AIDA stack and
bPlast and realigned the cables so that they
 11x 
  723   Fri Oct 24 10:31:59 2025 JB, GB, MP, AM, NH207Bi source threshold tests> > 
> > We have dismounted the AIDA stack and
bPlast and realigned the cables so that they
 6x 
  722   Thu Oct 23 18:45:12 2025 TDThursday 23 October 2025
19.45 Remote system check
 6x 
ELOG V3.1.3-7933898